OpenStudy (anonymous):

Find four consecutive integers whose sum is 114.

OpenStudy (anonymous):

set up an equation like this n+n+1+n+2+n+3=114

OpenStudy (anonymous):

If integers are consecutive, then they're in the form: x, x + 1, x + 2.... Since we got 4 consecutive integers, and we assign the value 'x' to the first time, we get : x, x + 1, x + 2, x + 3 We know the sum of these 4 terms is 114, so: (x) + (x + 1) + (x + 2) + (x + 3) = 114 4x + 6 = 114 4x = 108 x = 27 So now our 4 terms become: 27, 27 + 1, 27 + 2, 27 + 3, which is 27, 28, 29, 30

geerky42 (geerky42):